# FAST Polarization Catalog of FRB 20240114A

### 21 Mar 2026 v1

@@ V.3 Comparison with Other Repeaters and the Origin of Periodicity
The possible quasi-periodic modulation of burst activity in FRB 20240114A reported by Uttarkar et al. (2026), introduces an additional layer of complexity to the understanding of this source. Using Murriyang observations, they proposed a timescale of approximately 70 days and interpreted it as evidence for plasma lensing due to extreme scattering events. Zhou et al. (2025b) have also identified periodicities in FAST observations of this source, with periods of days (5.2 significance) and days (3.3 significance). Notably, periodic phenomena are not unique to FRB 20240114A. Several well-studied repeating FRBs have been reported to exhibit periodicities in either their burst activity or environmental properties, including FRB 20180916B (16.35-day activity period; Chime/Frb Collaboration et al., 2020), FRB 20220529 (200-day RM periodicity; Liang et al., 2025), and FRB 20201124A (26.24-day RM periodicity; Xu et al., 2025). The diversity of timescales and observational manifestations ranging from burst activity modulation to coherent RM variations suggests that multiple physical mechanisms may be at play. The interpretation of any potential periodicity in FRB 20240114A thus remains an open question, necessitating further monitoring and multi-wavelength observations to distinguish between intrinsic and environmental origins.
@@ VI Summary and Data Access
In this paper, we have presented the first large-scale polarization catalog for the repeating FRB 20240114A, resulting from an extensive monitoring campaign with the FAST telescope between 2024 January 28 and 2025 May 30. The core of this work is the detection and parameterization of 6,131 bursts. We have detailed the comprehensive data processing pipeline, from initial data acquisition and pre-processing to sophisticated polarization calibration and RM synthesis. The key deliverables of this pipeline are precise measurements for each burst, including the topocentric arrival time (MJD), DM, pulse width, bandwidth (BW), RM, and the degrees of linear and circular polarization.
@@ Appendix A Polarization Calibration
Saturation occurs in the FAST receiver when radio flux densities reach levels between hundreds of Jansky and mega-Jansky. To identify potential saturation events and mitigate their impact on polarization analysis, each burst’s four polarization components (XX, YY, X∗Y, Y∗X) were systematically examined. Saturation manifests differently across components: XX and YY saturate when intensities exceed 255 (wrapping to 0), while X∗Y and Y∗X saturate when absolute intensities surpass 127 (also wrapping to 0). For every burst, the dynamic spectra of all four polarization components were inspected. Critical-value data points (corresponding to the 255 and 127 thresholds) were flagged on the total intensity (I) dynamic spectra. These points were then frequency-integrated to generate time profiles. Two distinct scenarios were identified: Time profiles showing uniform temporal distribution indicate noise or radio frequency interference (RFI) effects (see Fig. 9). Time profiles matching the burst temporal structure signify genuine saturation (see Fig. 9). Bursts exhibiting the latter characteristic were excluded from subsequent analysis. It should be noted that the value 255 or 127 itself does not directly indicate saturation. The key criterion for identifying saturation is observing whether the distribution of tracers aligns with the pulse morphology. However, this serves only as an operational guideline, as the underlying principles are quite complex. Consequently, in practice there is no single unique method for determining saturation. Applying this methodology to the sample of bursts with SNR20 revealed 253 saturated events, which were subsequently removed to ensure reliable polarization measurements.
@@ Appendix B RM Synthesis
within a RM range of -20000 to 20000 , where . Due to challenges in calculating RM from weak signals or bursts with narrow emission bandwidths, RM values were successfully measured only 6,131 of the detected bursts. The mean value of the RM is 359.52 , showing significant variations throughout the observing campaign, ranging from 212.50 to 420.50 . We corrected the Stokes component and by derotating it using the measured RM. The degree of linear and circular polarization are calculated as,
@@ Appendix D Energy Distance
Note. In the main text, we employ the energy distance as a core metric to quantify the discrepancy between the linear and circular polarization fraction distributions of cumulative bursts in consecutive observing sessions with different time-scales.

### 18 Apr 2026 v2

@@ V.3 Comparison with Other Repeaters and the Origin of Periodicity
The possible quasi-periodic modulation of burst activity in FRB 20240114A reported by Uttarkar et al. (2026b), introduces an additional layer of complexity to the understanding of this source. Using Murriyang observations, they proposed a timescale of approximately 70 days and interpreted it as evidence for plasma lensing due to extreme scattering events. Zhou et al. (2025b) have also identified periodicities in FAST observations of this source, with periods of days (5.2 significance) and days (3.3 significance). Notably, periodic phenomena are not unique to FRB 20240114A. Several well-studied repeating FRBs have been reported to exhibit periodicities in either their burst activity or environmental properties, including FRB 20180916B, which shows a 16.35-day activity period (CHIME/FRB Collaboration et al., 2020). In this context, any potential activity periodicity in FRB 20240114A would place it among a small but growing group of repeaters with long-term activity modulation. The diversity of reported activity-modulation timescales among repeaters suggests that multiple physical mechanisms may be at play. The interpretation of any potential periodicity in FRB 20240114A thus remains an open question, necessitating further monitoring and multi-wavelength observations to distinguish between intrinsic and environmental origins.
@@ VI Summary and Data Access
In this paper, we present a large-scale polarization catalog for the repeating FRB 20240114A, resulting from an extensive monitoring campaign with the FAST telescope between 2024 January 28 and 2025 May 30. The core of this work is the detection and parameterization of 6,131 bursts. We have detailed the comprehensive data processing pipeline, from initial data acquisition and pre-processing to sophisticated polarization calibration and RM synthesis. The key deliverables of this pipeline are precise measurements for each burst, including the topocentric arrival time (MJD), DM, pulse width, bandwidth (BW), RM, and the degrees of linear and circular polarization.
@@ Appendix A Polarization Calibration
Saturation occurs in the FAST receiver when radio flux densities reach levels between hundreds of Jansky and mega-Jansky. To identify potential saturation events and mitigate their impact on polarization analysis, each burst’s four polarization components (XX, YY, X∗Y, Y∗X) were systematically examined. Here the quoted thresholds refer to the raw stored intensity values in the FAST FITS files rather than calibrated physical units. Saturation manifests differently across components: XX and YY saturate when intensities exceed 255 (wrapping to 0), while X∗Y and Y∗X saturate when absolute intensities surpass 127 (also wrapping to 0). For every burst, the dynamic spectra of all four polarization components were inspected. Critical-value data points (corresponding to the 255 and 127 thresholds) were flagged on the total intensity (I) dynamic spectra. These points were then frequency-integrated to generate time profiles. Two distinct scenarios were identified: Time profiles showing uniform temporal distribution indicate noise or radio frequency interference (RFI) effects (see Fig. 9). Time profiles matching the burst temporal structure signify genuine saturation (see Fig. 9). Bursts exhibiting the latter characteristic were excluded from subsequent analysis. It should be noted that the value 255 or 127 itself does not directly indicate saturation. The key criterion for identifying saturation is observing whether the distribution of tracers aligns with the pulse morphology. However, this serves only as an operational guideline, as the underlying principles are quite complex. Consequently, in practice there is no single unique method for determining saturation. Applying this methodology to the sample of bursts with SNR20 revealed 253 saturated events, which were subsequently removed to ensure reliable polarization measurements.
@@ Appendix B RM Synthesis
where . In our implementation, the trial RM range is set to -2000 to +2000 rad m-2. This interval was chosen as a conservative search range, substantially broader than the RMs measured for FRB 20240114A, while avoiding an unnecessarily wide search domain. The intrachannel depolarization factor can be written as , where in the narrow-channel approximation (Burn, 1966; Brentjens and de Bruyn, 2005). For the present FAST L-band channelization (0.122 MHz), 10% depolarization at the bottom of the band occurs only at rad m-2, far above both our adopted search range and the RMs measured in this work. The RM uncertainty is estimated empirically from the width of the RM-synthesis peak. Specifically, after identifying the RM corresponding to the maximum of the RM spectrum, we determine the left and right peak widths using the points where the spectrum drops below the midpoint between its maximum and minimum values, and divide these widths by the burst S/N. This yields asymmetric RM uncertainties when the RM-synthesis peak is not perfectly symmetric. For the present catalog, the RM analysis is restricted to bursts with S/N 20, for which the RM-synthesis peak is in most cases visually well defined. In practice, the corresponding polarization and RM-synthesis diagnostic plots were also inspected to identify obviously spurious cases. A more formal calibration of peak significance could in principle be obtained through dedicated Monte Carlo simulations, but this is beyond the scope of the present catalog paper. In addition to the empirical peak-width-based RM uncertainty quoted here, one RM-search grid step may be regarded as an additional systematic uncertainty ( 0.2 rad m-2) associated with the discretization of the RM search. Due to challenges in calculating RM from weak signals or bursts with narrow emission bandwidths, RM values were successfully measured only 6,131 of the detected bursts. The mean value of the RM is 359.52 , showing significant variations throughout the observing campaign, ranging from 212.50 to 420.50 . We corrected the Stokes component and by derotating it using the measured RM. The degree of linear and circular polarization are calculated as,
@@
For a subset of low-S/N bursts, the formally measured can exceed 100% (see Figure 5). These cases are preferentially associated with bursts whose intrinsic linear polarization is already very high, i.e. close to complete linear polarization. In such events, off-pulse noise fluctuations can occasionally yield in the background, and when the burst itself has relatively low integrated S/N, the accumulated result can produce . We therefore interpret these values not as physically meaningful super-100% polarization fractions, but as noise-biased measurements of bursts whose true is likely close to 100%.
@@ Appendix D Energy Distance
Note. In the main text, we use the energy distance to quantify the difference between the joint distributions of the linear and circular polarization fractions of cumulative bursts in consecutive observing sessions over different time-scales. Specifically, each burst is represented as a two-dimensional vector,
where and denote the linear and circular polarization fractions, respectively. For two samples and from two observing sessions, we compute the sample energy distance as
where is the Euclidean distance in the two-dimensional polarization-fraction space. A larger value of indicates a greater discrepancy between the two joint distributions.